# SETA: Semantic-Aware Token Augmentation for Domain Generalization

### v1

@@ IV-E Analytical Experiments
Ablation study on each component. We conduct the ablation study to investigate the efficacy of Energy-based Edge Token Selection (ETS) and Shape-disrupted Token Shuffling (STS) modules. Specifically, “Both Origin” is to mix representations of original sample with randomly selected sample representations. For “Edge + Origin”, we extract edge tokens of original sample and mix them with a randomly selected sample without token shuffling. “Origin + Shuffle” involves mixing the original sample representations with the token-shuffled sample representations. The experiments are conducted on the DomainBed benchmark using the GFNet-H-Ti backbone. As depicted in Tab. II, both the Energy-based Edge Tokens Selection (ETS) and the Shape-disrupted Token Shuffling (STS) consistently achieve improvements across five datasets, e.g., the “Edge + Origin” with ETS outperforms the “Both Origin” configuration by $0.41\%$ ($67.72\%$ vs. $67.31\%$), and the “Origin + Shuffle” with STS outperforms the “Both Origin” by $0.60\%$ ($67.91\%$ vs. $67.31\%$) on average. Furthermore, our SETA framework (i.e., “Edge + Shuffle”) achieves the best average performance, surpassing the Baseline by $1.62\%$ ($68.21\%$ vs. $66.59\%$). These results affirm that the ETS and STS modules complement and mutually enhance each other.
TABLE IV: Comparison (%) between the “Token Filtering” strategy and the “Fourier Masking” strategy. ”Token Filtering” involves filtering out edge-irrelevant tokens from feature maps, whereas ”Fourier Masking” refers to leveraging the high-frequency spectrum of feature maps for token mixing. The experiments are conducted on PACS using GFNet-H-Ti.
@@
Why filtering edge-irrelated tokens. To investigate the effect of the energy-based edge tokens selection (ETS) strategy, we here compare it with the “Fourier Masking” strategy that directly low-frequency filtering representation for token mixing. We here denote the ETS strategy as “Token Filtering”, which preserves a subset of original tokens that contain edge information. The experiment are conducted on the PACS dataset with the GFNet-H-Ti backbone. As presented in Tab. IV, the “Token Filtering” strategy achieves stable improvements from the “Fourier Masking” strategy, outperforming by $0.39\%$ ($90.52\%$ vs. $90.13\%$) on average. The results demonstrate that our ETS strategy could effectively protects useful semantic information within edge-related tokens, thus helping the model learn domain-invariant features.
Effects of SETA inserted positions. To assess the impact of the positions where SETA is inserted, given that a standard hierarchical model (for both GFNet-H-Ti and Swin-T) comprises four transformer layers as L1111-L4444, we evaluate the effects of applying SETA to different positions accordingly. As shown in Fig. 4 (a), each individually inserted SETA results in improvements compared to the baseline, with the best performance achieved when SETA is applied to all layers. The results suggest that augmenting shape bias in all network layers can effectively boost the generalization ability of the model. Therefore, we integrate SETA into all transformer layers for our experiments.
Sensitivity to edge token selection ratio. We here investigate the sensitivity of SETA to hyperparameter $p$ that denotes the percentage of edge tokens extracted from original samples. We vary the edge tokens percentage $p$ in $\{0.25,0.4,0.5,0.67,0.75,0.9,1.0\}$ on the PACS dataset. As presented in Fig. 4 (b), with $p$ increasing from $0.25$ to $0.75$, the accuracy slides from $88.28\%$ to $90.52\%$, which consistently exceeds the baseline by a large margin. The results prove the stability of our method for enhancing the shape bias of the model. However, is $p$ is excessively large, i.e., retaining too much texture tokens, the remaining shape-irrelevant features could cause the model to learn domain-specific information, resulting in a performance degradation of the model. The performance achieves the best when setting $p$ as $0.75$, which is adopted as default in all our experiments.
@@
Figure 4: Effects of hyper-parameters including inserted positions and edge token selection ratio in SETA. The experiments are conducted on PACS with GFNet-H-Ti as backbone. L1111-$4$ are four transformer layers of the network.
TABLE V: Fractions (%) of Dimensionality that encodes shape cues or texture cues [46, 23]. The texture and shape factors are estimated for the last-stage representation with the total dimension as $512$. The remaining dimensions are allocated to ”residual” factor.
@@
SETA (Ours) 33.83 (normal-$\uparrow$ 2.60) 18.67 (normal-$\downarrow$ 0.57)
@@
SETA-S-DSU (Ours) 34.31 (normal-$\uparrow$ 0.88) 18.54 (normal-$\downarrow$ 0.28)
@@
SETA-S-ALOFT (Ours) 34.15 (normal-$\uparrow$ 0.90) 18.56 (normal-$\downarrow$ 0.21)
Effectiveness on other architectures. To further validate the effectiveness of our method, we conduct experiments on various network architectures. We first provide the results of our SETA on the representative ViT model, i.e., DeiT-S [67] with $22$M parameters. For implementation details, we divide the $12$ network layers of DeiT-S into four blocks by $\{2,2,6,2\}$ as [18, 59], with SETA inserted into each block employing the same hyperparameters as previous experiments. The experiments are conducted on the PACS dataset. As presented in Tab. VI, on the advanced DeiT-S network, SETA still significantly enhances the model performance, achieving the improvement of $3.28\%$ ($89.13\%$ vs. $85.85\%$) from the baseline. The results highlight the effectiveness of our SETA method to enhance the shape bias of the model.
Moreover, while our SETA is a token-level augmentation designed for ViTs, we also design a variant for CNNs, denoted as SETA-CNN. Since CNNs utilize local convolutions for feature extraction, lacking patch-level token learning, the SETA-CNN is essentially a pixel-level feature augmentation. In detail, SETA-CNN treats each spatial pixel of feature maps in CNNs as a “token”, selecting edge-related pixels based on their values in edge maps, shuffling spatial pixel positions to generate shape-disrupted noise, and mixing them to produce pixel-level augmented representations. SETA-CNN is inserted after all residual blocks of the network. The experiments are conducted on the conventional ResNet-$50$ and the results are shown in Tab. VI. Our SETA achieves the improvement of $1.68\%$ ($87.18\%$ vs. $85.50\%$) from the baseline, which validates the generalizability of our method. However, CNNs primarily rely on local convolution for learning and lack exploration of spatial dependencies, while our method primarily enhances the shape bias of the model through feature extraction within tokens and spatial dependencies among tokens. The difference in architecture may constrain the performance of SETA on CNNs. In conclusion, the experimental results demonstrate the generalizability and effectiveness of our method.
TABLE VI: Performance (%) of SETA on othere architectures, including ResNet50505050 and DeiT. The experiment is conducted on the PACS dataset.
@@
Quantitative analysis of shape bias. To quantify the amount of texture and shape information, we follow the pioneering work [46] and estimate the number of channels that represent shape or texture concept. Given a representation $z\in\mathbb{R}^{N\times C}$, where $N$ is the token number and $C$ is the channel number. We initially construct image pairs $\{(z_{a},z_{b})\}$ with similar semantic concept $F$ (shapes or textures), by Fourier-based transformations [37], i.e., shape-bias pairs with the same phase but different amplitudes, and texture-bias pairs with the same amplitude but different phases. We approximate the mutual information between $z_{a}$ and $z_{b}$ with their correlation for each channel, summing over which yields a relative score for concept $F$, formulated as below:
@@
where $s_{F}$ is in $[-1,1]$. The remaining dimensions are allocated to residual factor $s_{r}$ with the maximum score $1$. For the set of concept factors $\{s_{r},s_{s},s_{t}\}$, we conduct a softmax to get the dimensionality fraction for the $F$ concept: $\rho_{F}=e^{s_{F}}/\sum_{k=1}^{3}e^{s_{k}}$. As shown in Tab. V, both the SOTA DG methods present a higher shape factor and lower texture factor than the baseline. The improvements are attributed to their perturbation of styles, potentially amplifying underlying shape bias. However, since they do not explicitly distinguish between global shape and local edges, their improvement to shape bias is still limited. Differently, our SETA perturbs local edges to enhance model learning of global shapes, thus effectively increasing shape factor while reducing texture factor, e.g., boosting shape factor by $0.88\%$ ($34.31\%$ vs. $33.43\%$) from DSU and $0.90\%$ ($34.15\%$ vs. $33.25\%$) from ALOFT. The results align well with the discussion in Sec. III-D, proving that our SETA can effectively enhance the shape sensitivity, thus improving the generalization ability of the model.
TABLE VII: The distribution gap between source and target domains. The experiments are conducted on the PACS dataset with GFNet-H-Ti as the backbone. The lower the value, the smaller the distribution gap.
@@
Domain gap between source and target domains. To investigate the impact of SETA on the domain gap between source and target domains, we estimate the domain distance $d_{H\Delta H}(D_{s}^{aug},D_{t})$ in Eq. (15) by evaluating the domain discrepancy of the feature maps extracted by the last layer of the model. Let $\overline{f_{s}^{k}}$ denote the averaged feature maps of all samples from the $k$-th source domain, and $\overline{f_{t}}$ represents the averaged feature maps of all samples from the target domain. The domain discrepancy between the source and target domains is then computed as follows:
$d=\frac{1}{K}\sum_{k=1}^{K}\|\overline{f_{s}^{k}}-\overline{f_{t}}\|_{2}.$ (25)
The experiments are conducted on the PACS dataset using the GFNet-H-Ti architecture, in which we select a domain as the target domain and evaluate the distance between it and the other source domains. As shown in Tab. VII, both DSU and ALOFT demonstrate the ability to reduce the domain gap between source and target domains, indicating that perturbing style features could enhance the model robustness to domain shifts. However, these methods overlook the fact that edge features could also encode domain-specific information (e.g., shape-irrelevant local edges), leading to an insufficient narrowing of the domain gap. In contrast, our proposed method explicitly encourages the model to learn the holistic shape of the object by perturbing local edges. Consequently, our method effectively reduces the domain gap between source and target domains, e.g., achieving a reduction of $4.43$ ($5.52$ vs. $9.95$) compared to the Baseline. Notably, with the target domain being Sketch emphasizing shape features, our SETA achieves a significant reduction in domain distance between Sketch and the other three source domains, reducing $8.73$ ($7.15$ vs. $15.88$) from the Baseline. The result indicates that our method could effectively enhance the model learning of shape representation. Combined with existing augmentation methods, our SETA further reduces domain gaps, thereby improving the model generalization. The results empirically validate the theoretical conclusions outlined in Sec. III-D.
TABLE VIII: The computational overhead of SETA and its two variants. The experiments are conducted on the PACS dataset.
Methods Training Time (s) GPU Mem (GB)
@@
Baseline [59] 0.220 9.827
SETA 0.226 9.867
DSU [14] 0.235 11.207
SETA-S-DSU 0.236 11.829
ALOFT [13] 0.242 12.026
SETA-S-ALOFT 0.246 12.537
@@
Baseline [18] 0.153 8.009
SETA 0.154 8.227
DSU [14] 0.181 8.245
SETA-S-DSU 0.184 8.467
ALOFT [13] 0.197 8.366
SETA-S-ALOFT 0.200 8.947
@@
Computational overhead of SETA. To showcase the computational efficiency of SETA, we conduct a comparative analysis including training time per iteration and GPU memory requirements. We compare our methods with the baseline [67] and SOTA augmentation methods in DG, i.e., DSU [14] and ALOFT [13]. The results are shown in Tab. VIII. Specifically, for SETA, we observe a marginal increase in both training time and GPU memory requirement, e.g., on Swin-T architecture, there is an incremental time of $0.006$ seconds per iteration ($0.226$ seconds vs. $0.220$ seconds), accompanied by an additional memory requirement of $0.040$ GB ($9.867$ GB vs. $9.827$ GB) during training. Moreover, when SETA is synergistically employed with existing DG augmentation methods, our variants still exhibit a negligible computational overhead, proving the efficiency of SETA and its two variants.
@@
Visualization on the synthetic changes. We here provide intuitive visualizations showcasing the diverse changes generated by our SETA method. Following [14, 77], we employ a pre-defined autoencoder, i.e., the VGG-$19$ network, to visualize the reconstructed results. Specifically, SETA is incorporated into the encoder, and the augmented representations are then inverted into synthetic images by the decoder. The visualization results are illustrated in Fig. 5. Comparing the reconstructed images obtained with and without SETA, it is evident that our method introduces diverse variations by perturbing both the style information and local edges while preserving the shape structures. Consequently, with the integration of SETA, shape-irrelevant features are perturbed, compelling the model to learn domain-invariant global information and incentivizing shape bias of the model.
@@
Figure 6: Visualization of attention maps in each layer if the network on the PACS dataset with the Art Painting as the target domain. The backbone used in the experiment is the GFNet-H-Ti architecture.
@@
Visualization of attention maps. To visually verify the claim that the SETA can enhance the model’s learning of global shape information, we present the attention maps of each network layer for baseline and SETA methods, employing the GradCAM technique [78]. As presented in Fig. 6, the baseline model predominantly directs its attention towards numerous shape-irrelevant regions across all network layers, which mainly encode domain-related information, thereby leading the model to overfit the source domains. In contrast, the SETA effectively perturbs the shape-irrelevant features at each layer, compelling the model to learn a more comprehensive understanding of global shape representations. Taking the giraffe for example, distinct features such as its face and neck, which are pivotal for classification, are accurately captured by the SETA. Conversely, the baseline model tends to focus on background and texture features, thus leadeing to misclassification. These results suggest that the SETA can effectively incentivize the shape bias of the model and facilitate the learning of generalized structural information.

### v2

@@ IV-E Analytical Experiments
TABLE III: Ablation study (%) on different modules of SETA on the DomainBed benchmark. We denote the Activation-based Edge Tokens Selection as ETS and the Shape-disrupted Token Shuffling as STS, respectively. The experiments are conducted on GFNet-H-Ti.
Methods ETS STS PACS VLCS OfficeHome TerraInc DomainNet Avg.
TABLE IV: Comparison (%) between the “Token Filtering” strategy and the “Fourier Masking” strategy. “Token Filtering” involves filtering out edge-irrelevant tokens from feature maps, whereas “Fourier Masking” refers to leveraging the high-frequency spectrum of feature maps for token mixing. The experiments are conducted on PACS using GFNet-H-Ti.
@@
Ablation study on each component. We conduct the ablation study to investigate the efficacy of Activation-based Edge Token Selection (ETS) and Shape-disrupted Token Shuffling (STS) modules. Specifically, “Both Origin” is to mix representations of the original sample with randomly selected sample representations. For “Edge + Origin”, we extract edge tokens of the original sample and mix them with a randomly selected sample without token shuffling. “Origin + Shuffle” involves mixing the original sample representations with the token-shuffled sample representations. The experiments are conducted on the DomainBed benchmark using the GFNet-H-Ti backbone. As depicted in Tab. III, both ETS and STS consistently achieve improvements across five datasets, e.g., the “Edge + Origin” with ETS outperforms the “Both Origin” configuration by $0.41\%$ ($67.72\%$ vs. $67.31\%$), and the “Origin + Shuffle” with STS outperforms the “Both Origin” by $0.60\%$ ($67.91\%$ vs. $67.31\%$) on average. Furthermore, our SETA framework (i.e., “Edge + Shuffle”) achieves the best average performance, surpassing the Baseline by $1.62\%$ ($68.21\%$ vs. $66.59\%$). These results affirm that the ETS and STS modules complement and mutually enhance each other.
TABLE V: Effects of different weights $w$ for edge-irrelated tokens in our SETA. The experiment is conducted on the PACS dataset with GFNet-H-Ti as the backbone.
Weight $w$ Art Cartoon Photo Sketch Avg.
1.00 90.86 85.26 98.34 85.35 89.95
0.75 90.16 84.51 98.28 86.16 89.78
0.50 90.64 85.11 98.13 86.19 90.02
0.25 90.48 85.30 98.28 86.23 90.07
0.00 90.95 85.55 98.37 87.19 90.52
Why filtering edge-irrelated tokens. To investigate the effect of the activation-based edge tokens selection (ETS) strategy, we here compare it with the “Fourier Masking” strategy that directly low-frequency filtering representation for token mixing. We here denote the ETS strategy as “Token Filtering”, which preserves a subset of original tokens that contain edge information. The experiments are conducted on the PACS dataset with the GFNet-H-Ti backbone. As presented in Tab. IV, the “Token Filtering” strategy achieves stable improvements from the “Fourier Masking” strategy, outperforming by $0.39\%$ ($90.52\%$ vs. $90.13\%$) on average. The results demonstrate that our ETS strategy could effectively protect useful semantic information within edge-related tokens, thus helping the model learn domain-invariant features.
To study whether simply masking the edge token affects feature extraction, we also conduct an experiment that leverages weights to weight non-edge tokens instead of roughly masking them. Specifically, we introduce a hyperparameter $w$ to assign weights to non-edge tokens, where the Eq. (5) is re-formulated as: $z_{i}^{e}=\mathcal{M}_{edge}\odot z_{i}+w\cdot(1-\mathcal{M}_{edge})\odot z_{i}$. We investigate the impact of $w$ on model generalization, where $w\in\{0.00,0.25,0.50,0.75,1.00\}$. Notably, $w=0.00$ corresponds to the non-edge masking method (i.e., SETA), while $w=1.00$ is the baseline without SETA. As presented in Tab. V, SETA (i.e., $w=0.00$) achieves the best performance among all variants, e.g., exceeding the variant with $w=0.75$ by $0.74\%$ ($90.52\%$ vs. $89.78\%$). The results indicate that SETA could adaptively retain tokens that contain most of the edge information while introducing the non-edge tokens with weight $w$ could introduce harmful domain-specific information.
@@
Figure 5: Effects of hyper-parameters including inserted positions and low-frequency mask scale $r$ in SETA. The experiments are conducted on PACS with GFNet-H-Ti backbone. L1111-$4$ are four transformer layers of the network.
(a) On PACS.
(b) On OfficeHome.
Figure 6: Effects of the hyper-parameter edge token selection ratio $p$ in SETA. The experiments are conducted on PACS and OfficeHome with GFNet-H-Ti.
Effects of SETA inserted positions. To assess the impact of the positions where SETA is inserted, given that a standard hierarchical model (for both GFNet-H-Ti and Swin-T) comprises four transformer layers as L1111-L4444, we evaluate the effects of applying SETA to different positions accordingly. As shown in Fig. 5 (a), each individually inserted SETA results in improvements compared to the baseline, with the best performance achieved when SETA is applied to all layers. The results suggest that augmenting shape bias in all network layers can effectively boost the generalization ability of the model. Therefore, we integrate SETA into all transformer layers for all our experiments.
Sensitivity to low-frequency mask scale. We analyze the effect of the hyperparameter $r$, which determines the extent to which low-frequency components are discarded, on the generalization ability of the model. We vary $r$ within the range $\{0.2,0.4,0.6,0.8,1.0\}$ on PACS. As shown in Fig. 5 (b), the best performance is achieved with $r=0.6$ for both datasets, suggesting that this setting effectively filters out low-frequency style information while preserving high-frequency edge information. Thus, we use $r=0.6$ across all datasets.
Sensitivity to edge token selection ratio. We here investigate the sensitivity of SETA to the hyperparameter $p$, which represents the percentage of edge tokens extracted from the original samples. We vary $p$ within the range $\{0.25,0.4,0.5,0.67,0.75,0.9,1.0\}$ on both PACS and OfficeHome. As shown in Fig. 6, as $p$ increases from $0.25$ to $0.75$, the accuracy improves from $88.28\%$ to $90.52\%$ on PACS and from $71.66\%$ to $72.33\%$ on OfficeHome, consistently surpassing the baseline by a significant margin. These results prove the stability of our method in enhancing the shape bias of the model. However, if $p$ is too large, i.e., retaining too many texture tokens, the remaining shape-irrelevant features may cause the model to learn domain-specific information, leading to performance degradation. The optimal performance is achieved with $p$ set to $0.75$, which we adopt as the default setting in our experiments for all datasets.
TABLE VI: Fractions (%) of Dimensionality that encodes shape cues or texture cues [43, 21]. The texture and shape factors are estimated for the last-stage representation with the total dimension as $512$. The remaining dimensions are allocated to the “residual” factor.
@@
SETA (Ours) 33.83 ($\uparrow$ 2.60) 18.67 ($\downarrow$ 0.57)
@@
SETA-S-DSU (Ours) 34.31 ($\uparrow$ 0.88) 18.54 ($\downarrow$ 0.28)
@@
SETA-S-ALOFT (Ours) 34.15 ($\uparrow$ 0.90) 18.56 ($\downarrow$ 0.21)
Effectiveness on other architectures. To further validate the effectiveness of our method, we conduct experiments on various network architectures. We first provide the results of our SETA on the representative ViT model, i.e., DeiT-S [64] with $22$M parameters. For implementation details, we divide the $12$ network layers of DeiT-S into four blocks by $\{2,2,6,2\}$ as [18, 56], with SETA inserted into each block employing the same hyperparameters as previous experiments. The experiments are conducted on the PACS dataset. As presented in Tab. VII, on the advanced DeiT-S network, SETA still significantly enhances the model performance, achieving the improvement of $3.28\%$ ($89.13\%$ vs. $85.85\%$) from the baseline. The results highlight the effectiveness of our SETA method to enhance the shape bias of the model.
Moreover, while our SETA is a token-level augmentation designed for ViTs, we also design a variant for CNNs, denoted as SETA-CNN. Since CNNs utilize local convolutions for feature extraction, lacking patch-level token learning, the SETA-CNN is essentially a pixel-level feature augmentation. In detail, SETA-CNN treats each spatial pixel of feature maps in CNNs as a “token”, selecting edge-related pixels based on their values in edge maps, shuffling spatial pixel positions to generate shape-disrupted noise, and mixing them to produce pixel-level augmented representations. SETA-CNN is inserted after all residual blocks of the network. The experiments are conducted on the conventional ResNet-$50$, and the results are shown in Tab. VII. Our SETA achieves the improvement of $1.68\%$ ($87.18\%$ vs. $85.50\%$) from the baseline, which validates the generalizability of our method. However, CNNs primarily rely on local convolution for learning and lack exploration of spatial dependencies, while our method primarily enhances the shape bias of the model through feature extraction within tokens and spatial dependencies among tokens. The difference in architecture may constrain the performance of SETA on CNNs. In conclusion, the experimental results demonstrate the generalizability and effectiveness of our method.
TABLE VII: Performance (%) of SETA on othere architectures, including ResNet50505050 and DeiT. The experiment is conducted on the PACS dataset.
@@
Quantitative analysis of shape bias. To quantify the amount of texture and shape information, we follow the pioneering work [43] and estimate the number of channels that represent the shape or texture concept. Given a representation $z\in\mathbb{R}^{N\times C}$, where $N$ is the token number and $C$ is the channel number. We initially construct image pairs $\{(z_{a},z_{b})\}$ with similar semantic concept $F$ (shapes or textures), by Fourier-based transformations [34], i.e., shape-bias pairs with the same phase but different amplitudes, and texture-bias pairs with the same amplitude but different phases. We approximate the mutual information between $z_{a}$ and $z_{b}$ with their correlation for each channel, summing over which yields a relative score for concept $F$, formulated as below:
@@
where $s_{F}$ is in $[-1,1]$. The remaining dimensions are allocated to residual factor $s_{r}$ with the maximum score $1$. For the set of concept factors $\{s_{r},s_{s},s_{t}\}$, we conduct a softmax to get the dimensionality fraction for the $F$ concept: $\rho_{F}=e^{s_{F}}/\sum_{k=1}^{3}e^{s_{k}}$. As shown in Tab. VI, both the SOTA DG methods present a higher shape factor and lower texture factor than the baseline. The improvements are attributed to their perturbation of styles, potentially amplifying underlying shape bias. However, since they do not explicitly distinguish between global shape and local edges, their improvement to shape bias is still limited. Differently, our SETA perturbs local edges to enhance model learning of global shapes, thus effectively increasing shape factor while reducing texture factor, e.g., boosting shape factor by $0.88\%$ ($34.31\%$ vs. $33.43\%$) from DSU and $0.90\%$ ($34.15\%$ vs. $33.25\%$) from ALOFT. The results align well with the discussion in Sec. III-D, proving that our SETA can effectively enhance the shape sensitivity, thus improving the generalization ability of the model.
TABLE VIII: The distribution gap between source and target domains. The experiments are conducted on the PACS dataset with GFNet-H-Ti as the backbone. The lower the value, the smaller the distribution gap.
@@
Domain gap between source and target domains. To investigate the impact of SETA on the domain gap between source and target domains, we estimate the domain distance $d_{H\Delta H}(D_{s}^{aug},D_{t})$ in Eq. (15) by evaluating the domain discrepancy of the feature maps extracted by the last layer of the model. Let $\overline{f_{s}^{k}}$ denote the averaged feature maps of all samples from the $k$-th source domain, and $\overline{f_{t}}$ represents the averaged feature maps of all samples from the target domain. The domain discrepancy between the source and target domains is then computed as: $d=\frac{1}{K}\sum_{k=1}^{K}\|\overline{f_{s}^{k}}-\overline{f_{t}}\|_{2}.$ The experiments are conducted on the PACS dataset using the GFNet-H-Ti architecture, in which we select a domain as the target domain and evaluate the distance between it and the other source domains. As shown in Tab. VIII, both DSU and ALOFT demonstrate the ability to reduce the domain gap between source and target domains, indicating that perturbing style features could enhance the model robustness to domain shifts. However, these methods overlook the fact that edge features could also encode domain-specific information (e.g., shape-irrelevant local edges), leading to an insufficient narrowing of the domain gap. In contrast, our proposed method explicitly encourages the model to learn the holistic shape of the object by perturbing local edges. Consequently, our method effectively reduces the domain gap between source and target domains, e.g., achieving a reduction of $4.43$ ($5.52$ vs. $9.95$) compared to the Baseline. Notably, with the target domain being Sketch emphasizing shape features, our SETA achieves a significant reduction in domain distance between Sketch and the other three source domains, reducing $8.73$ ($7.15$ vs. $15.88$) from the Baseline. The result indicates that our method could effectively enhance the model learning of shape representation. Combined with existing augmentation methods, our SETA further reduces domain gaps, thereby improving the model generalization. The results empirically validate the theoretical conclusions outlined in Sec. III-D.
TABLE IX: The computational overhead of SETA and its two variants. The experiments are conducted on the PACS dataset.
Methods Training Time (s) GPU Mem (GB) Performance (%)
@@
Baseline [56] 0.220 9.827 66.92
SETA 0.226 10.328 67.50 ($\uparrow 0.58$)
DSU [14] 0.235 11.207 67.29
SETA-S-DSU 0.236 11.829 67.81 ($\uparrow 0.52$)
ALOFT [13] 0.242 12.026 67.38
SETA-S-ALOFT 0.246 12.537 68.00 ($\uparrow 0.62$)
@@
Baseline [18] 0.153 8.009 66.59
SETA 0.154 8.227 68.21 ($\uparrow 1.62$)
DSU [14] 0.181 8.245 67.33
SETA-S-DSU 0.184 8.467 68.36 ($\uparrow 1.03$)
ALOFT [13] 0.197 8.366 68.04
SETA-S-ALOFT 0.200 8.947 68.58 ($\uparrow 0.54$)
@@
Computational overhead of SETA. To showcase the computational efficiency of SETA, we conduct a comparative analysis including training time per iteration and GPU memory requirements. We compare our methods with the baseline [64] and SOTA augmentation methods in DG, i.e., DSU [14] and ALOFT [13]. The results are shown in Tab. IX. Specifically, for SETA, we observe a marginal increase in both training time and GPU memory requirement, e.g., on Swin-T architecture, there is an incremental time of $0.006$ seconds per iteration ($0.226$ seconds vs. $0.220$ seconds), accompanied by an additional memory requirement of $0.040$ GB ($9.867$ GB vs. $9.827$ GB) during training. Moreover, when SETA is synergistically employed with existing DG augmentation methods, our variants still exhibit a negligible computational overhead, proving the efficiency of SETA and its two variants.
@@
Visualization on the synthetic changes. We here provide intuitive visualizations showcasing the diverse changes generated by our SETA method. Following [14, 75], we employ a pre-defined autoencoder, i.e., the VGG-$19$ network, to visualize the reconstructed results. Specifically, SETA is incorporated into the encoder, and the augmented representations are then inverted into synthetic images by the decoder. The visualization results are illustrated in Fig. 7. Comparing the reconstructed images obtained with and without SETA, it is evident that our method introduces diverse variations by perturbing both the style information and local edges while preserving the shape structures. Consequently, with the integration of SETA, shape-irrelevant features are perturbed, compelling the model to learn domain-invariant global information and incentivizing the shape bias of the model.
@@
Figure 8: Visualization of attention maps in each layer of the network on the PACS dataset with Art Painting as the target domain. The backbone used in the experiment is the GFNet-H-Ti architecture.
@@
Visualization of attention maps. To visually verify the claim that the SETA can enhance the model’s learning of global shape information, we present the attention maps of each network layer for baseline and SETA methods, employing the GradCAM technique [76]. As presented in Fig. 8, the baseline model predominantly directs its attention towards numerous shape-irrelevant regions across all network layers, which mainly encode domain-related information, thereby leading the model to overfit the source domains. In contrast, the SETA effectively perturbs the shape-irrelevant features at each layer, compelling the model to learn a more comprehensive understanding of global shape representations. Taking the giraffe for example, distinct features such as its face and neck, which are pivotal for classification, are accurately captured by the SETA. Conversely, the baseline model tends to focus on background and texture features, thus leading to misclassification. These results suggest that the SETA can effectively incentivize the shape bias of the model and facilitate the learning of generalized structural information.